OpenStudy (anonymous):

I tried my calculator and won't give me the right answer. 11.98 100.00 23.85 Add 5% tax 10% discount What is the total?

jaynator495 (jaynator495):

128.36 is what i got Explnation 11.98+100+23.85 A=? 5% of A=? (make sure to remove the extra decimal points threw this problem) Add 5% of A to current total 10% of total=? subtract 10% of total to get your final awnser Final Answer=128.36 you can fill in the blanks hope this helped -Jaynator495

jaynator495 (jaynator495):

man that took a while to type xD

jaynator495 (jaynator495):

sorry about that taking so long to type lol

OpenStudy (anonymous):

128.36 is it

jaynator495 (jaynator495):

see... i was right xP

jaynator495 (jaynator495):

@AquaStone316

jaynator495 (jaynator495):

@Hotchellerae21 @cj49 @ChloBeiiar15 Would you agree with my solution guys?

OpenStudy (anonymous):

@Jaynator495 i agree

OpenStudy (anonymous):

11.98+100.00+23.85=135.83(.90)=122.247+(122.247*.05)=128.36

jaynator495 (jaynator495):

cody you didnt mention that you converted the % into decimal and thats what .90 and .05 is, it doesnt help if he doesnt know what it means you know?

OpenStudy (anonymous):

Thank you, by the way I'm a girl.

jaynator495 (jaynator495):

ok ._.